OpenStudy (anonymous):

Write an equation of the tangent function with period 3pi/8, phase shift -pi/5, vertical shift -2.

OpenStudy (noelgreco):

What have you tried? Where do you fail to understand the question?

OpenStudy (anonymous):

I wasn't at school and missed the whole unit. Was given no notes, no textbook for the class. Just need help knowing what to do. Even an example similar to this one I could probably solve it.

OpenStudy (aum):

If the tangent function is: y = tan(Bx + C) + D then: Period = pi / B Phase Shift = -C / B Vertical shift = D

OpenStudy (aum):

You can find B, C and D from: "period 3pi/8, phase shift -pi/5, vertical shift -2"

OpenStudy (anonymous):

That's not one of my choices though. I got something different

OpenStudy (aum):

Just noticed my mistake: period = pi / B = 3pi/8 divide both sides by pi: 1/B = 3/8 cross multiply: 8 = 3B B = 8/3 (had the reciprocal before) phase shift = - C/B = -pi/5 multiply both sides by -1: C/B = pi/5 cross multiply: 5C = pi * B = pi * 8/3 divide by 5: C = pi * 8/3 * 1/5 = 8pi/15 D = vertical shift = -2 Therefore, the tan function is tan(Bx + C) + D = tan(8/3*x + 8*pi/15) - 2